# A halo model approach for mock catalogs of time-variable strong gravitational lenses

### v4

@@ Abstract
Time delays in both galaxy- and cluster-scale strong gravitational lenses have recently attracted a lot of attention in the context of the Hubble tension. Future wide-field cadenced surveys, such as the Legacy Survey of Space and Time (LSST), are anticipated to discover strong lenses across various scales. We generate mock catalogs of strongly lensed quasars (QSOs) and supernovae (SNe) on galaxy-, group-, and cluster-scales based on a halo model that incorporates dark matter halos, galaxies, and subhalos. For the upcoming LSST survey, we predict that approximately 3500 lensed QSOs and 200 lensed SNe with resolved multiple images will be discovered. Among these, about 80 lensed QSOs and 10 lensed SNe will have maximum image separations larger than $10~{}\mathrm{arcsec}$, which roughly correspond to cluster-scale strong lensing. We find that adopting the Chabrier stellar initial mass function (IMF) instead of the fiducial Salpeter IMF reduces the predicted number of strong lenses approximately by half, while the distributions of lens and source redshifts and image separations are not significantly changed. In addition to mock catalogs of multiple-image lens systems, we create mock catalogs of highly magnified systems, including both multiple-image and single-image systems. We find that such highly magnified systems are typically produced by massive galaxies, but non-negligible fraction of them are located in the outskirt of galaxy groups and clusters. Furthermore, we compare subsamples of our mock catalogs with lensed QSO samples constructed from the Sloan Digital Sky Survey and Gaia to find that our mock catalogs with the fiducial Salpeter IMF reproduce the observation quite well. In contrast, our mock catalogs with the Chabrier IMF predict a significantly smaller number of lensed QSOs compared with observations, which adds evidence that the stellar IMF of massive galaxies is Salpeter-like. Our python code to generate the mock catalogs, Strong Lensing Halo model-based mock catalogs (SL-Hammocks), as well as mock catalogs of lensed QSOs and SNe are made available online.
@@ 1 Introduction
In recent years, the field of cosmology has been confronted with a significant challenge known as the Hubble tension problem (see, e.g., Di Valentino et al., 2021, for a review). The Hubble tension is an empirical problem in which two types of observations have conflicting measurements of the Hubble constant $H_{0}$. In the first approach, $H_{0}$ is estimated from the early-universe measurements, in particular cosmic microwave background (CMB), assuming the so-called standard cosmology, $\Lambda$-dominated cold dark matter ($\Lambda$CDM) model. For example, Planck Collaboration (2020) provides a tight constraint of $H_{0}=67.4\pm 0.5\mathrm{~{}km}\mathrm{~{}s}^{-1}\mathrm{Mpc}^{-1}$ assuming the flat $\Lambda$CDM model. The other approach measures $H_{0}$ more directly from observations of the local Universe. Recently, Supernovae and $H_{0}$ for the Equation of State of dark energy (SH0ES) collaboration has obtained $H_{0}=73.04\pm 1.04\mathrm{~{}km}\mathrm{~{}s}^{-1}\mathrm{Mpc}^{-1}$ by type Ia supernovae (SNe Ia) calibrated via the distance ladder in the local Universe (Riess et al., 2022). The two types of observations exhibit inconsistency with significant $5\sigma$ tension of the value of $H_{0}$, which may pose a problem for the validity of the $\Lambda$CDM model.
@@
One of the important avenues for resolving this problem is the estimation of the Hubble constant with other independent methodologies. The Tip of the Red Giant Branch (TRGB), a waypoint in the evolutionary state for giant stars, offers another way to determine $H_{0}$ as a local standard candle. Some work (see e.g., Freedman et al., 2020) shows a consistent result with both to $1.5\sigma$ or less of $H_{0}\sim 70~{}\mathrm{~{}km}\mathrm{~{}s}^{-1}\mathrm{Mpc}^{-1}$, while others (see e.g., Anand et al., 2020; Blakeslee et al., 2021; Jones et al., 2022) yield somewhat higher value at $H_{0}\sim 71.5-73~{}\mathrm{~{}km}\mathrm{~{}s}^{-1}\mathrm{Mpc}^{-1}$. Recently, Scolnic et al. (2023) measure $H_{0}=73.22\pm 2.06\mathrm{~{}km}\mathrm{~{}s}^{-1}\mathrm{Mpc}^{-1}$ using TRGB, which further clarifies the tension, by analyzing in conjunction with the Pantheon$+$ SN Ia sample while considering a correction for differences in SN surveys and local flows. Various other methods have also been used to measure $H_{0}$, including Megamasers (Kuo et al., 2013; Reid et al., 2013; Kuo et al., 2015; Reid et al., 2019), gravitational waves (Abbott et al., 2017; Mooley et al., 2018; Hotokezaka et al., 2019), fast radio bursts (Hagstotz et al., 2022; Wu et al., 2022; James et al., 2022; Liu et al., 2023a; Zhao et al., 2022), baryon acoustic oscillations (Addison et al., 2018; Benisty and Staicova, 2021), Type II SNe (de Jaeger et al., 2022), ages of old astrophysical objects (Vagnozzi et al., 2022; Wei and Melia, 2022), and others (Moresco et al., 2022). However, a satisfactory answer to this tension has not been found yet.
@@
Another independent approach for $H_{0}$ measurements is to focus on time delays between multiple images of strong gravitational lenses. It was pointed out in the 1960s that observations of time delays between multiple images of gravitationally lensed SNe can be used to estimate $H_{0}$ independently of the distance ladder (Refsdal, 1964). However, it is only recently that we have been able to measure the Hubble constant accurately enough to use it to potentially resolve the Hubble tension.
@@
Due to the scarcity of lensed SNe, lensed quasars (QSOs) have been mostly used to constrain the Hubble constant. For example, in Wong et al. (2020), the $H_{0}$ Lenses in COSMOGRAIL’s Wellspring (H0LiCOW) collaboration estimated $H_{0}$ as $H_{0}=73.3_{-1.8}^{+1.7}\mathrm{~{}km}\mathrm{~{}s}^{-1}\mathrm{Mpc}^{-1}$ based on analysis of the time delays of six lensed QSOs with the simple mass distribution of lensing galaxies such as an elliptical power-law or stars plus standard dark matter halos. In Birrer et al. (2020), the time-delay cosmography (TDCOSMO) team obtained a constraint of $H_{0}=74.5_{-6.1}^{+5.6}\mathrm{~{}km}\mathrm{~{}s}^{-1}\mathrm{Mpc}^{-1}$, by considering a more flexible radial mass distribution and breaking the degeneracy in mass models with stellar kinematics. While galaxy-scale lensed QSOs are used in the H0LiCOW and TDCOSMO analyses, recently, in Kelly et al. (2023), SN Refsdal, the first multiply-lensed SN with time delay measurements, is finally used to obtain the constraint on the Hubble constant of $H_{0}=64.8_{-4.3}^{+4.4}\mathrm{~{}km}\mathrm{~{}s}^{-1}\mathrm{Mpc}^{-1}$ (see also Liu and Oguri, 2024, for the detailed analysis of the mass model dependence). In addition to the difference in the source population, SN Refsdal was strongly lensed by a massive cluster of galaxies rather than a single galaxy and, hence, is highly complementary to the constraints from H0LiCOW and TDCOSMO. Besides, another measurement of $H_{0}$ from the gravitationally lensed SN H0pe due to a massive cluster of galaxies, $H_{0}=75.7_{-5.5}^{+8.1}\mathrm{~{}km}\mathrm{~{}s}^{-1}\mathrm{Mpc}^{-1}$, was reported very recently (Pascale et al., 2024).
@@
The measurement of $H_{0}$ using time-delay lensings will advance significantly in the coming decade thanks to several large-scale galaxy survey projects, including Euclid and Rubin Observatory Legacy Survey of Space and Time (LSST). Oguri and Marshall (2010, OM10 hereafter) was among the first to make a comprehensive prediction of event rates of galaxy-scale lensed SNe and QSOs, such as those used in the H0LiCOW and TDCOSMOS analysis, in various wide-field cadenced imaging surveys including LSST. According to OM10, the LSST should find more than 8000 galaxy-scale lensed QSOs, some 3000 of which will have well-measured time delays. The LSST should also find some 130 lensed SNe during the 10-year survey duration. OM10 also computed the available precision on the Hubble constant from the LSST survey and obtained $\sigma(h)=0.017$ as the predicted marginalized $68$ percent confidence intervals, where $h\equiv H_{0}/100~{}\mathrm{km}\mathrm{~{}s}^{-1}\mathrm{Mpc}^{-1}$. We note that there are some minor updates in the calculation of OM10, as presented in Oguri (2018); Lemon et al. (2023); Shen et al. (2023). Henceforth, we refer to the updated OM10 calculation as OM10+.
@@
Strong gravitational lenses are predominantly produced by single massive galaxies. This is one of the reasons why the analysis of TDCOSMO and H0LiCOW, as well as OM10, have focused on galaxy-scale strong lenses. However, several cluster-scale lensed QSO systems have already been found, which recently attracts increasing attention due to the Hubble tension problem. Denzel et al. (2021) used the cluster-scale lensed QSO SDSS J1004+4112 (Inada et al., 2003; Oguri et al., 2004; Inada et al., 2005; Fohlmeister et al., 2008) together with 7 galaxy-scale lensed QSOs to obtain $H_{0}=71.8_{-3.3}^{+3.9}\mathrm{~{}km}\mathrm{~{}s}^{-1}\mathrm{Mpc}^{-1}$. Napier et al. (2023) focused more specifically on cluster-scale lensed QSOs by studying three cluster-lensed QSOs SDSS J1004+4112, SDSS J1029+2623 (Inada et al., 2006; Oguri et al., 2008a; Fohlmeister et al., 2013), and SDSS J2222+2745 (Dahle et al., 2013, 2015), to produce a combined measurement of the Hubble constant as $H_{0}=74.1\pm 8.0\mathrm{~{}km}\mathrm{~{}s}^{-1}\mathrm{Mpc}^{-1}$. Very recently, SDSS J1004+4112 has been reanalyzed by Martínez-Arrizabalaga et al. (2023) using a free-form lens model and by Liu et al. (2023b) using 16 different lens models, with results of $H_{0}=74_{-13}^{+9}\mathrm{~{}km}\mathrm{~{}s}^{-1}\mathrm{Mpc}^{-1}$ and $H_{0}=67.5_{-8.9}^{+14.5}\mathrm{~{}km}\mathrm{~{}s}^{-1}\mathrm{Mpc}^{-1}$, respectively.
@@
In this work, we present detailed predictions of the numbers of gravitationally lensed QSOs and SNe produced for a wide mass range of lens objects, including both galaxy- and cluster-scale lenses, with a particular focus on the LSST. Specifically, we adopt the so-called halo model (see e.g., Asgari et al., 2023, for a review) approach adopting a compound lens model that consists of the dark matter halo and galaxy (stellar mass) components. This approach differs from the one in OM10 and OM10+ that focuses only on galaxy-scale strong lenses. They assume the singular isothermal ellipsoid (SIE) model and calculate the lensing properties from the line-of-sight projected stellar velocity dispersion of lens galaxies obtained from the distributions of the velocity dispersions, the so-called velocity function, empirically derived in the local universe by the Sloan Digital Sky Survey (SDSS). While the SIMCT pipeline presented in More et al. (2016) can also generate mock lenses at galaxy- as well as group- and cluster-scales, their approach differs from the halo model approach in this paper in that they use a hybrid approach of injecting mock lensed sources in foreground deflectors selected from a real or mock survey.
@@
In addition to mock catalogs of multiply lens systems as those presented in OM10+, we construct mock catalogs of highly magnified systems, for which both single and multiple image systems are included. Since the contribution from group-scale halos is expected to be important for predicting the abundance of highly magnified systems (Keeton et al., 2005), we expect that our halo model approach enables a reliable prediction on the abundance and properties of such systems. The mock catalogs of highly magnified systems will be useful for e.g., assessing the magnification bias on observations of high-redshift QSOs and SNe in the LSST survey.
@@
Throughout this paper, we assume a flat $\Lambda$CDM cosmology and fix the cosmological parameters to the Planck 2018 best-fits (Planck Collaboration, 2020), specifically $h=0.677$, $\Omega_{\mathrm{b}}h^{2}=0.0224,\omega_{\mathrm{cdm}}=0.120,\ln\left(10^{10}\mathcal{A}_{\zeta}\right)=3.05$, and $n_{\mathrm{s}}=0.967$.
@@ 2 Deflector model
This section addresses the halo-model-based deflector model, which distinguishes our new mock catalog from that of OM10. In this model, we generate deflector populations from dark matter halos and subhalos. We employ a compound mass model where the mass distribution of individual deflectors is represented by a superposition of dark matter and stellar components. The advantage of this compound model over the simple SIE model used in OM10 is that it can naturally account for the transition in the density profile from the galaxy- to the cluster-mass scale, reflecting the structure formation through dark matter and baryon (e.g., Kochanek and White, 2001; Oguri, 2006; More et al., 2012). While OM10 provides the distribution of image separations for galaxy-scale lenses expected to be found in the LSST, this study extends the scope to include not only galaxy-scale lenses but also group- and cluster-scale lenses. Since early-type galaxies, particularly elliptical galaxies, are known to account for approximately $80\%$ of the total lensing probability (Turner et al., 1984; Fukugita et al., 1992; Kochanek, 1996; Möller et al., 2007), we focus solely on elliptical galaxies as the stellar components of deflectors throughout this paper.

### v5

@@ Abstract
Time delays in both galaxy- and cluster-scale strong gravitational lenses have recently attracted a lot of attention in the context of the Hubble tension. Future wide-field cadenced surveys, such as the Legacy Survey of Space and Time (LSST), are anticipated to discover strong lenses across various scales. We generate mock catalogs of strongly lensed quasars (QSOs) and supernovae (SNe) on galaxy-, group-, and cluster-scales based on a halo model that incorporates dark matter halos, galaxies, and subhalos. For the upcoming LSST survey, we predict that approximately 4000 lensed QSOs and 200 lensed SNe with resolved multiple images will be discovered. Among these, about 80 lensed QSOs and 10 lensed SNe will have maximum image separations larger than $10~\mathrm{arcsec}$, which roughly correspond to cluster-scale strong lensing. We find that adopting the Chabrier stellar initial mass function (IMF) instead of the fiducial Salpeter IMF reduces the predicted number of strong lenses approximately by half, while the distributions of lens and source redshifts and image separations are not significantly changed. In addition to mock catalogs of multiple-image lens systems, we create mock catalogs of highly magnified systems, including both multiple-image and single-image systems. We find that such highly magnified systems are typically produced by massive galaxies, but non-negligible fraction of them are located in the outskirt of galaxy groups and clusters. Furthermore, we compare subsamples of our mock catalogs with lensed QSO samples constructed from the Sloan Digital Sky Survey and Gaia to find that our mock catalogs with the fiducial Salpeter IMF reproduce the observation quite well. In contrast, our mock catalogs with the Chabrier IMF predict a significantly smaller number of lensed QSOs compared with observations, which adds evidence that the stellar IMF of massive galaxies is Salpeter-like. Our python code to generate the mock catalogs, Strong Lensing Halo model-based mock catalogs (SL-Hammocks), as well as mock catalogs of lensed QSOs and SNe are made available online.
@@ 1 Introduction
[…]
@@
One of the important avenues for resolving this problem is the estimation of the Hubble constant with other independent methodologies. The Tip of the Red Giant Branch (TRGB), a waypoint in the evolutionary state for giant stars, offers another way to determine $H_{0}$ as a local standard candle. Some work (see e.g., 2020ApJ...891...57F) shows a consistent result with both to $1.5\sigma$ or less of $H_{0}\sim 70~\mathrm{~km}\mathrm{~s}^{-1}\mathrm{Mpc}^{-1}$, while others (see e.g., 2020AAS...23633601A; 2021ApJ...911...65B; 2022ApJ...933..172J) yield somewhat higher value at $H_{0}\sim 71.5-73~\mathrm{~km}\mathrm{~s}^{-1}\mathrm{Mpc}^{-1}$. Recently, 2023ApJ...954L..31S measure $H_{0}=73.22\pm 2.06\mathrm{~km}\mathrm{~s}^{-1}\mathrm{Mpc}^{-1}$ using TRGB, which further clarifies the tension, by analyzing in conjunction with the Pantheon++ SN Ia sample while considering a correction for differences in SN surveys and local flows. Various other methods have also been used to measure $H_{0}$, including Megamasers (2013ApJ...767..155K; 2013ApJ...767..154R; 2015ApJ...800...26K; 2019ApJ...886L..27R), gravitational waves (2017Natur.551...85A; 2018Natur.561..355M; 2019NatAs...3..940H), fast radio bursts (2022MNRAS.511..662H; 2022MNRAS.515L...1W; 2022MNRAS.516.4862J; 2023ApJ...946L..49L; 2022arXiv221213433Z), baryon acoustic oscillations (2018ApJ...853..119A; 2021A&A...647A..38B), Type II SNe (2022MNRAS.514.4620D), ages of old astrophysical objects (2022JHEAp..36...27V; 2022ApJ...928..165W), and others (2022LRR....25....6M). However, a satisfactory answer to this tension has not been found yet.
@@
Another independent approach for $H_{0}$ measurements is to focus on time delays between multiple images of strong gravitational lenses. It was pointed out in the 1960s that observations of time delays between multiple images of gravitationally lensed SNe can be used to estimate $H_{0}$ independently of the distance ladder (1964MNRAS.128..307R). However, it is only recently that we have been able to measure the Hubble constant accurately enough to use it to potentially resolve the Hubble tension.
@@
Due to the scarcity of lensed SNe, lensed quasars (QSOs) have been mostly used to constrain the Hubble constant. For example, in 2020MNRAS.498.1420W, the $H_{0}$ Lenses in COSMOGRAIL’s Wellspring (H0LiCOW) collaboration estimated $H_{0}$ as $H_{0}=73.3_{-1.8}^{+1.7}\mathrm{~km}\mathrm{~s}^{-1}\mathrm{Mpc}^{-1}$ based on analysis of the time delays of six lensed QSOs with the simple mass distribution of lensing galaxies such as an elliptical power-law or stars plus standard dark matter halos. In 2020A&A...643A.165B, the time-delay cosmography (TDCOSMO) team obtained a constraint of $H_{0}=74.5_{-6.1}^{+5.6}\mathrm{~km}\mathrm{~s}^{-1}\mathrm{Mpc}^{-1}$, by considering a more flexible radial mass distribution and breaking the degeneracy in mass models with stellar kinematics. While galaxy-scale lensed QSOs are used in the H0LiCOW and TDCOSMO analyses, recently, in 2023Sci...380.1322K, SN Refsdal, the first multiply-lensed SN with time delay measurements, is finally used to obtain the constraint on the Hubble constant of $H_{0}=64.8_{-4.3}^{+4.4}\mathrm{~km}\mathrm{~s}^{-1}\mathrm{Mpc}^{-1}$ (see also 2024arXiv240213476L, for the detailed analysis of the mass model dependence). In addition to the difference in the source population, SN Refsdal was strongly lensed by a massive cluster of galaxies rather than a single galaxy and, hence, is highly complementary to the constraints from H0LiCOW and TDCOSMO. Besides, another measurement of $H_{0}$ from the gravitationally lensed SN H0pe due to a massive cluster of galaxies, $H_{0}=75.7_{-5.5}^{+8.1}\mathrm{~km}\mathrm{~s}^{-1}\mathrm{Mpc}^{-1}$, was reported very recently (2024arXiv240318902P).
@@
The measurement of $H_{0}$ using time-delay lensings will advance significantly in the coming decade thanks to several large-scale galaxy survey projects, including Euclid and Rubin Observatory Legacy Survey of Space and Time (LSST). 2010MNRAS.405.2579O was among the first to make a comprehensive prediction of event rates of galaxy-scale lensed SNe and QSOs, such as those used in the H0LiCOW and TDCOSMOS analysis, in various wide-field cadenced imaging surveys including LSST. According to 2010MNRAS.405.2579O, the LSST should find more than 8000 galaxy-scale lensed QSOs, some 3000 of which will have well-measured time delays. The LSST should also find some 130 lensed SNe during the 10-year survey duration. 2010MNRAS.405.2579O also computed the available precision on the Hubble constant from the LSST survey and obtained $\sigma(h)=0.017$ as the predicted marginalized $68$ percent confidence intervals, where $h\equiv H_{0}/100~\mathrm{km}\mathrm{~s}^{-1}\mathrm{Mpc}^{-1}$. We note that there are some minor updates in the calculation of 2010MNRAS.405.2579O, as presented in 2018MNRAS.480.3842O; 2023MNRAS.520.3305L; 2023ApJ...943...38S. Henceforth, we refer to the updated 2010MNRAS.405.2579O calculation as OM10+.
@@
Strong gravitational lenses are predominantly produced by single massive galaxies. This is one of the reasons why the analysis of TDCOSMO and H0LiCOW, as well as 2010MNRAS.405.2579O, have focused on galaxy-scale strong lenses. However, several cluster-scale lensed QSO systems have already been found, which recently attracts increasing attention due to the Hubble tension problem. 2021MNRAS.501..784D used the cluster-scale lensed QSO SDSS J1004+4112 (2003Natur.426..810I; 2004ApJ...605...78O; 2005PASJ...57L...7I; 2008ApJ...676..761F) together with 7 galaxy-scale lensed QSOs to obtain $H_{0}=71.8_{-3.3}^{+3.9}\mathrm{~km}\mathrm{~s}^{-1}\mathrm{Mpc}^{-1}$. 2023ApJ...959..134N focused more specifically on cluster-scale lensed QSOs by studying three cluster-lensed QSOs SDSS J1004+4112, SDSS J1029+2623 (2006ApJ...653L..97I; 2008ApJ...676L...1O; 2013ApJ...764..186F), and SDSS J2222+2745 (2013ApJ...773..146D; 2015ApJ...813...67D), to produce a combined measurement of the Hubble constant as $H_{0}=74.1\pm 8.0\mathrm{~km}\mathrm{~s}^{-1}\mathrm{Mpc}^{-1}$. Very recently, SDSS J1004+4112 has been reanalyzed by 2023arXiv230914776M using a free-form lens model and by 2023PhRvD.108h3532L using 16 different lens models, with results of $H_{0}=74_{-13}^{+9}\mathrm{~km}\mathrm{~s}^{-1}\mathrm{Mpc}^{-1}$ and $H_{0}=67.5_{-8.9}^{+14.5}\mathrm{~km}\mathrm{~s}^{-1}\mathrm{Mpc}^{-1}$, respectively.
@@
In this work, we present detailed predictions of the numbers of gravitationally lensed QSOs and SNe produced for a wide mass range of lens objects, including both galaxy- and cluster-scale lenses, with a particular focus on the LSST. Specifically, we adopt the so-called halo model (see e.g., 2023OJAp....6E..39A, for a review) approach adopting a compound lens model that consists of the dark matter halo and galaxy (stellar mass) components. This approach differs from the one in 2010MNRAS.405.2579O and OM10+ that focuses only on galaxy-scale strong lenses. They assume the singular isothermal ellipsoid (SIE) model and calculate the lensing properties from the line-of-sight projected stellar velocity dispersion of lens galaxies obtained from the distributions of the velocity dispersions, the so-called velocity function, empirically derived in the local universe by the Sloan Digital Sky Survey (SDSS). While the SIMCT pipeline presented in 2016MNRAS.455.1191M can also generate mock lenses at galaxy- as well as group- and cluster-scales, their approach differs from the halo model approach in this paper in that they use a hybrid approach of injecting mock lensed sources in foreground deflectors selected from a real or mock survey.
@@
In addition to mock catalogs of multiply lens systems as those presented in OM10+, we construct mock catalogs of highly magnified systems, for which both single and multiple image systems are included. Since the contribution from group-scale halos is expected to be important for predicting the abundance of highly magnified systems (2005ApJ...621..559K), we expect that our halo model approach enables a reliable prediction on the abundance and properties of such systems. The mock catalogs of highly magnified systems will be useful for e.g., assessing the magnification bias on observations of high-redshift QSOs and SNe in the LSST survey.
@@
[…]
@@ 2 Deflector model
This section addresses the halo-model-based deflector model, which distinguishes our new mock catalog from that of 2010MNRAS.405.2579O. In this model, we generate deflector populations from dark matter halos and subhalos. We employ a compound mass model where the mass distribution of individual deflectors is represented by a superposition of dark matter and stellar components. The advantage of this compound model over the simple SIE model used in 2010MNRAS.405.2579O is that it can naturally account for the transition in the density profile from the galaxy- to the cluster-mass scale, reflecting the structure formation through dark matter and baryon (e.g., 2001ApJ...559..531K; 2006MNRAS.367.1241O; 2012ApJ...749...38M). While 2010MNRAS.405.2579O provides the distribution of image separations for galaxy-scale lenses expected to be found in the LSST, this study extends the scope to include not only galaxy-scale lenses but also group- and cluster-scale lenses. Since early-type galaxies, particularly elliptical galaxies, are known to account for approximately $80\%$ of the total lensing probability (1984ApJ...284....1T; 1992ApJ...393....3F; 1996ApJ...466..638K; 2007MNRAS.379.1195M), we focus solely on elliptical galaxies as the stellar components of deflectors throughout this paper.